Accessibility Assistance

Skip to Content




Information Technology
George A. Smathers Libraries
University of Florida
Suite 1250
2124 NE Waldo Rd
Gainesville, FL 32609

Telephone: (352) 273-2800
Fax: (352) 392-2354

Information Technology Staff Meeting Notes
October 5, 2010

Present: Shruti Bagri, Thomas Bielicke, Cynthia Bowker, Laurie Brennan, Will Chaney, Logan Clapp, Matthew Daley, Debra Fetzer, Winston Harris, Michael Jay, Maureen Kelly, Tom Minton, Chris Nicolich, Amy Polk, Josh Spurgin, Evan Wack
Announcements
  1. Open Enrollment - Sept 27-Oct 22.
  2. Winston thanked Cynthia & Debra for their work on the IT Open House.
Ongoing IT Updates

Hardware
  1. McAfee - EPO is still not reporting. Trish is on the phone right now with McAfee - recreating filters.
  2. Altiris - holding off on configuration until the server move is completed.
  3. SharePoint hardware site is growing. Once we have the templates, then we'll be ready to launch. There really hasn't been any progress on the templates. General discussion on the SharePoint site. It was decided that we will have a meeting with all involved staff to figure out the direction of the IT SharePoint site.
  4. Software re-organization - records in Grover are 80% wrong. Win7 staff migration on hold until we can get the software situation straightened out.
  5. Room 1A - podium fixed. They are working on the sound today.
  6. Win7 public - Clean Slate fix did not work. Hope to have it fixed and be able to migrate the public machines by the fall-spring intercession.
  7. Waiting for vault item pickup by Surplus Property. Need to survey old server items as well.
  8. Thanks to all for their help last week with the public machines that were removed from the domain.
Software
  1. OCR automation - Chris was researching the Kirtas solution last week. It will not work in the way it was envisioned to function. Mark is looking at just using ABBY. He is trying to get a free evaluation copy.
  2. Document Disposition database - demoed internally, looks really good. There are some changes to be made. We invited Jan Swanbeck, etc. out next week for another demo. She is going to a conference in Washington, may show a screen-shot demo of the database.
  3. Macros - completions, additions. 10 tickets currently open.
  4. OLE - meeting this morning. Week-long training on Kuali-Rice was highly informative. The project is moving along.
  5. We had one application that did not transfer to the new server. Bill Hanssen was unable to look up the PID field in Grover.
  6. 2-3 new projects: new candidate assessment, attendance counter for public library meetings and HR recruitment study.
Network/Server Administration
  1. SharePoint - Logan to schedule a meeting with all parties involved.
  2. App server - being copied. Prep work tomorrow. Increasing drives to an extra 4G. Will check permissions. Will bring it up Thursday morning.
  3. Web done today. Going to a new box tonight, then we'll bring the old box down.
  4. After that, only a few servers left: Illiad, NT18, NT15.
  5. Joe is ready to do the GIS install.
  6. MSL server is totally offline. Logan and Josh have moved it to the Eastside. Will install Altiris on it next week.
ILS
  1. Cynthia had two technical issues that brought her work to a halt: Donna (FCLA) asked her to do a task that involved installing Oracle and the other issue was that she did not have permissions to run ALEPH services.
DLC General Issues
  1. Waiting on the OCR issue.
  2. Finishing up documentation.
  3. Matt Loving wants a domain linked to DLC.
  4. OCR Prime is working.
Web Projects
  1. Oil spill site - no contact in a long while.
  2. Judaica - Rebecca contacted Matthew with some content, but it is not up yet.
  3. SharePoint page for document user orientation - training online.